Changes between Initial Version and Version 1 of UC20200803


Ignore:
Timestamp:
Jul 30, 2021, 7:49:19 AM (13 months ago)
Author:
trac
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • UC20200803

    v1 v1  
     1== APO 3.5-m Users Committee Telecon, 08/03/2020 ==
     2
     3Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Mike Skrutskie (UVa), Ben Williams (UW), Misty Bentz (GSU), Chip Kobulnicky (UWy), Denise Stephens (BYU), Russet !McMillan (APO), Mukremin Kilic (OU)
     4
     5----
     6
     7=== User feedback and comments from institutional representatives: ===
     8
     9* NMSU - nothing to report
     10* Washington - nothing to report
     11* Oklahoma - nothing to report
     12* NAPG - no report
     13* Colorado - nothing to report
     14* Wyoming - nothing to report
     15* Seattle - no report
     16* FGCU - no report
     17* UVa - nothing to report
     18* Georgia State - nothing to report
     19* JHU - nothing to report
     20* BYU - nothing to report
     21
     22----
     23
     24=== Telescope and Instruments Report: ===
     25
     26Jamey's detailed report is included below, followed by additional information discussed during today's call.
     27
     28''3.5-m Telescope and Instruments Highlights, 7/03/2020 - 7/29/2020''
     29
     300) Overview
     31
     32APO has been open for night operations since June 1. We are still restricting all visitors to the site. Following the state of New Mexico and NMSU guidelines, we also remain under travel restrictions ourselves.   
     33 
     34On-site staffing levels have been increased slightly from the minimal staff allowed on site before 6/1/20. Currently, as NMSU is in Phase 3 of the Return to Research, APO is allowed between 35-50% of staffing. The NMSU numbers, coupled with the state of New Mexico restrictions for a non-essential classified business like APO, which allow staffing levels of 50% of the Fire Marshall determined maximum occupancy of a building, has allowed us to have more people on site at any given time. 
     35 
     36Weather for the month of July has been as expected for monsoon rain season. We are losing time consistently since the beginning of the month due to weather. We are experiencing an above average year of Miller moths and as such are still dealing with them. Happily, the moth issue has abated considerably since the heavier monsoon rains have started. Keeping fingers crossed that the issue doesn’t get worse again in the fall.
     37
     381) Telescope
     39
     40The telescope has been performing as expected.  There was an issue with the tertiary rotation (post shutdown) that we believe has been fixed by enabling the correction factors for movement. 
     41 
     42Summer shutdown tasks were performed and the telescope was returned to operations on schedule. Pointing data were taken and a new model was put into operation. 
     43 
     44Recently (7/26) we believe that we may have experienced a lightning strike on/near the 3.5m enclosure. The telescope exhaust fan system appears to have some contactors that have one phase damaged. Replacement parts are on order.
     45
     462) Instruments
     47
     48DIS: Scattered light is still an issue with the instrument. It hasn’t been used much since the last report, so there is no new information on the previously mentioned turret (and/or detent) motor appearing to be acting up. The grating issue that was affecting blue, a mechanical issue, has been resolved. Recurrence of the issue is not expected.
     49
     50!TripleSpec: powered on, cold and operational.
     51
     52Agile: powered on, cold and operational.
     53
     54Echelle: powered on, cold and operational. The inter-order light (IOL) for red and blue continue to drop, but at a slower rate than last month.
     55
     56NIC-FPS: Frequent expose inits are being required to remedy an all-zero readout problem. The corrupted readouts are not new but seem to have changed their pattern, apparently coming in clusters sometimes and not at other times. Previous known triggers such as instrument handling or rotating the instrument upside-down do not seem to be as likely to trigger the issue now, but we haven't really figured out what - if anything - is a common trigger. Investigations are ongoing.
     57
     58ARCTIC: powered on, cold, and operational.
     59
     603) ARCSAT
     61
     62ARCSAT is operational and has not been used for many observations in the past month;  weather has been the main limiting factor. Some observing time was lost due to the telescope having the wrong time, causing a dome/telescope mismatch. A low flux level streaking issue in !SurveyCam images is being investigated.
     63
     64----
     65
     66=== Additional telescope and instrument discussion: ===
     67
     68We are now back to operating every night. The COVID restrictions that are in place are not impacting the nighttime operations as much as daytime because engineering daytime support (on-site) is currently more limited. Many thanks to Russet for making herself available for some daytime troubleshooting. Regarding the telescopes, we thought we had lightning strike that resulted in several system failures, although now we are not as convinced that was the cause. In any case, everything is now back to operating normally. The DIS scattered light is still an issue. Russet and Ed fixed the grating issue that was affecting the blue side. It was first thought to be an electronic issue and then mechanical; now it is working. There are frequent readout issues with NICFPS that will be investigated this week. It may be rotationally dependent. There have been limited observations with ARCSAT lately due to weather. We will be working with the !SurveyCam camera supplier to address the streaking issue.
     69
     70----
     71
     72=== ARCTIC Flats ===
     73
     74With the old NA2 baffle we were aware that dome flats were not as good as sky flats for ARCTIC, and they were only good if one was only concerned with the central quarter of the chip. Some of this was alleviated with the new baffle, which now provides an unvignetted field all the way out to edge of chip, but now we see a gradient across the chip when taking dome flats. As a reminder, when taking dome flats the lamps illuminate the mirror covers. We have recommended for quite some time that users should not take dome flats with Agile. We are still exploring this issue when it comes to ARCTIC flats, and we are seeking volunteers to help with more analysis. The new baffle allows ARCTIC to see the mirror covers when they are closed. This doesn’t happen when the mirror covers are open, therefore this doesn’t affect twilight (sky) flats. There is no official change in our recommendations yet. Separately, we are aware of ARCTIC flat irregularities reported by those using very blue filters, but we think that is an issue of a differential response. The dome flats look very different from sky flats in the bluest filters due to the difference in color between the source lamps and the sky. '''We recommend that users not use dome flats for ARCTIC when using those very blue filters (broad or narrowband), and more generally, try to use sky flats whenever possible.'''
     75
     76----
     77
     78=== APO and the COVID-19 virus ===
     79
     80We have additional protocols in place to allow for a limited number of contractors to be on site. Visitors coming from out-of-state have to self-quarantine for 14 days (per New Mexico law) before coming to APO. Currently we can handle 1-2 additional visitors, and we can house them at Sunspot if we have to. Jack is now allowed to return work on-site, which will alleviate the short-handedness the 3.5m observing specialists had been experiencing. They still have to do extra cleanings each night.
     81
     82----
     83=== Q3 3.5m scheduling: ===
     84
     85There is a little bit of OPEN time left in Q3. For people who want to request some of the short time slots, if they request to use the instrument that is already mounted on the telescope they could potentially share flats and skip focusing, which would save some time.
     86
     87'''UC reps: notify users that they should let Ben, Russet and Nancy know if they are interested in requesting any of the remaining OPEN/DD time. For an existing program this requires simply sending an email; if the request is not associated with an existing program already scheduled in Q3 they will ask for a short science justification.'''
     88----
     89
     90=== Q3 ARCSAT scheduling: ===
     91
     92There is some OPEN ARCSAT time in September.
     93
     94'''UC reps: have your users let Ben, Nancy and Russet know if they are interested in requesting any of the remaining ARCSAT time.'''
     95
     96----
     97
     98=== Q4 3.5m scheduling: ===
     99
     100The institutional schedulers have all received their Q4 allocations. Ben asked that all requests be submitted no later than August 26. There is no special time set aside for KOSMOS commissioning in Q4, but we are adding a separate training session each quarter (see below).  BYU will also have a training session for their users in Q4.
     101
     102----
     103
     104=== Updated Training Policies and Procedures: ===
     105
     106With the new COVID travel and site visitation restrictions still in place (and likely to remain in place for a while), we are trying to address how to effectively substitute the on-site certification training experience with a virtual one. Candace has put together a great training program that can be done online, but it actually takes MORE time to complete it than to do a training in person (although there is no travel time for the trainees). It requires more effort on Candace's part since it involves several Zoom meetings ahead of time, days/weeks prior to the actual training/observing. It is therefore ''critically'' important that anyone interested in getting trained arrange it well ahead of time. Class "trips" that are now being done virtually get scheduled in advance through the standard proposal submission process, so there should be no change for that procedure. However, sending a postdoc, student, or new faculty member to the site for a one-off training session can no longer be accommodated. To that end, we are implementing a single training session each quarter that individuals from ARC member institutions and leasing partners can sign up for.   This should ease the scheduling burden and make it more predictable for everyone. The dates for the Q4 training session are not yet known, but they will be published in the Q4 schedule when it comes out. Those interested in taking advantage of that opportunity in a given quarter should send an email (or have it sent by the UC rep or institutional scheduler) to Ben, Russet, Candace, and Nancy. We will evaluate this strategy after we see how it goes in Q4!
     107----
     108
     109=== KOSMOS Update: ===
     110
     111The TEG is allowed back in the lab on a limited basis, so Rishi has been in lab working on trying to track down the leak in the new dewar.
     112
     113----
     114
     115=== ARCTIC Diffuser Rotation ===
     116
     117Parts have been ordered and some bearings are being tested.
     118
     119----
     120
     121=== DAWG Update ===
     122
     123The DAWG met and is working to have a white paper completed by early September. It will identify priorities for the observatory in terms of data storage and software.
     124
     125----
     126
     127=== Science Communications Update ===
     128
     129Prior to the meeting Nancy sent around a link to a pilot "science hangout" done with four Astronomy graduate students from NMSU who are using APO for their dissertation research. The feedback was fairly positive, although the general feeling was that it was too long (~ 20 minutes). It was suggested that we make shorter videos, one for each person ~ 5-7 minutes long. There was some discussion of who the intended audience is (answer: there are two -- the general public, and astronomers at current/future ARC member institutions). For the videos for the public, rather than a monologue it might be better to have a Q & A format with pretty pictures and more basic explanations of what people are studying. Relatedly, it was suggested that we should host an online day-long science meeting of APO users so that everyone could learn about the other science being done with the 3.5m. This suggestion was well-received by the Users Committee reps, and the suggestion was that we plan it for some time in 2021 (spring?). Nancy also expressed an interest in working with Jamey on scheduling some "science hangouts" with the APO staff so that they could learn more about the science being done with the 3.5m if interested.
     130
     131=== ACTION ITEMS:===
     132
     133'''Open action items from previous meetings:'''
     134
     135* ARCTIC filter order - Nancy.  Status: OPEN. After hearing back from several people who expressed opinions about filter bandwidths (with no overwhelming consensus), she is ready to move forward, but has also received some names of additional vendors to contact. She will get some more price quotes and then place the order.
     136
     137* UC reps: notify users that we are operating with fewer obs-specs than normal and there is a chance this could impact our observing schedule.  Also, please ask users to be patient with the night staff, and to respond to the reminder emails in a timely fashion. Status: CLOSED.
     138
     139* UC reps: notify users that there is a limited amount of OPEN/DD time remaining on both the 3.5m and ARCSAT during Q3. If interested in requesting some time, please follow standard protocols for making such requests. Status: OPEN.
     140
     141* UC reps: instruct users that if they have strong feelings about whether we should go with very narrow or less narrow Ha etc. filters for ARCTIC, they should notify Nancy ASAP. Status: CLOSED.
     142
     143* UC reps: remind all users to send their list of refereed/submitted publications from 2019-present to Nancy by 7/31/2020. Status: CLOSED (but she will still take them!).
     144
     145'''New action items from this meeting:'''
     146
     147* UC reps: remind all users to use twilight flats with ARCTIC rather than dome flats when possible, and to consider contributing flats to the APO library of ARCTIC flats that is being built up (contact Russet !McMillan for details).
     148
     149* Science highlights are due to Nancy by October 15 (see All Other Business below).
     150
     151----
     152
     153=== All Other Business ===
     154
     155Nancy issued her annual call for science highlights from late 2019-present for presentation at the ARC Board of Governors meeting in early November.  They are due to Nancy no later than October 15, and can be collected by a single representative from each institution and mailed as a packet (easier for Nancy!) or sent to her directly by individual investigators. There is no slide template, but each slide should contain at minimum: project title, author(s) and affiliation(s), a graphic, a few text bullets describing the result, and a citation if there is one.
     156
     157----
     158
     159=== Next meeting:===
     160
     161The next meeting will be on Tuesday September 1 at 10:30 am MDT. NOTE: the meeting schedule was changed to be ALL TUESDAYS for the remainder of the calendar year. Thus the remaining meetings in 2020 will be on September 1, October 6, November 3, and December 8.
     162
     163----